For the given functions f(x)= 3/(x-4) and g(x) = x/2 find (f(g(x)). I started the problem by substituting x/2 in for f(x), so I get to f(g(x)) = 3/((x/2) -4).
Answer by stanbon(75887)

For the given functions f(x)= 3/(x-4) and g(x) = x/2 find (f(g(x)). I started the problem by substituting x/2 in for f(x), so I get to f(g(x)) = 3/((x/2) -4).
-------------------------
That is correct.
Now simplify:
f[g(x)] = 3/[(x-8)/2]
= 6/(x-8)
=============
Cheers,
Stan H.